OpenStudy (anonymous):

12 4 --- (D) --- (D is dividing) 17 17 A 48 B 17 C 4 D 3

OpenStudy (anonymous):

(12/17)/(4/17)=(12+17)/(17*4)=12/4=3 D

OpenStudy (anonymous):